# Rethinking Oversmoothing in Graph Neural Networks: A Rank-Based Perspective

### v3

@@ 6 Experiments
Empirical studies on the evolution of oversmoothing measures often use untrained, hundred-layer-deep GNNs [35, 44, 34, 45]. We emphasize that this is an overly simplified setting. In more realistic settings, as the ones considered in this section, a trained GNN may suffer from significant performance degradation after only few-layers, at which stage the convergence patterns of most of the metrics are difficult to observe. The experiments that we present in this section validate the robustness of the relaxed rank metrics in quantifying oversmoothing in GNNs against the other metrics.
@@
  • The Dirichlet Energy $E_{\mathrm{Dir}}$ [4, 33] and its variant $E_{\mathrm{Proj}}$ [45]. Both are discussed in Section 3.1, see in particular (1) and (2).
  • Normalized versions of Dirichlet energy and its variant, $E_{\mathrm{Dir}}(X)/\|X\|_{F}^{2}$ and $E_{\mathrm{Proj}}(X)/\|X\|_{F}$. Indeed, from our previous discussion, a robust oversmoothing measure should be scale invariant with respect to the features. Metrics with global normalization like the ones we consider here have also been proposed in [8, 30, 20].
  • The Mean Average Distance (MAD) [5]
@@
  • Relaxed rank metrics: We consider the Numerical Rank and Effective Rank. Both are discussed in Section 5. We point out that from our theoretical investigation, in particular from (4), the numerical rank decays to $1$ faster than the decay of the normalized $E_{\mathrm{Proj}}$ energy to zero. This further supports the use of the Numerical Rank as an improved measure of oversmoothing with respect to $E_{\mathrm{Proj}}$.
Figure 2: Four examples of the metric behaviours computed at the last hidden layer of separately trained GCNs of increasing depths. For Erank and Numrank, we measure Erank$(X)-r^{*}_{\mathrm{ER}}$ and $\mathrm{NumRank}(X)-r^{*}_{\mathrm{NR}}$ for some $r^{*}>1$. In these particular cases, $r^{*}_{\mathrm{ER}}<1.85$, $r^{*}_{\mathrm{NR}}<1.3$. Note that the effective rank and numerical rank of the input features $X^{(0)}$ are about 1084 and 13.6, respectively. Additional results are attached in Appendix G.
In Table 1 and Figure 2, we train GNNs of a fixed hidden dimension equal to 32 on homophilic, heterophilic and large-scale datasets in their default splits. We follow the standard setups of GCN and GAT as stated in Sections 2.1 and 2.2, and use homogeneous LeakyReLU (LReLU) as the activation function. For each configuration, GNNs of eight different depths ranging from 2 to 24 are trained. The oversmoothing metric and accuracy results are averaged over 10 separately trained GNNs. All GNNs are trained with NAdam Optimizer and a constant learning rate of 0.01. The oversmoothing metrics are computed at the last hidden layer before the output layer. In Figure 2 and in Appendix G, we plot the behaviour of the different oversmoothing measures, the norm of the features, and the accuracy of the trained GNNs with increasing depth. These figures clearly show that the network suffers a significant drop in accuracy, which is not matched by any visible change in standard oversmoothing metrics. By contrast, the rank of the feature representations decreases drastically, following closely the behaviour of the network’s accuracy. These findings are further sustained by the results shown in Table 1, where we compute the Pearson correlation coefficient between the logarithm of every measure and the classification accuracy of every GNN model. The use of a logarithmic transformation is based on the understanding that oversmoothing grows exponentially with the length of the network.
Extensions of these results are provided in Appendices E and F. In Appendix D, we perform an asymptotic ablation study on very deep (300-layer) synthetic networks with randomly sampled, untrained weights. This study serves to validate our theoretical findings on the convergence of relaxed rank metrics and to demonstrate that such untrained settings offer little insight into the ability of existing metrics to quantify oversmoothing in realistic, trained networks.
@@ 7 Conclusion
In this paper, we have discussed the problem of quantifying oversmoothing in message-passing GNNs. After simplifying the existing theoretical analysis using nonlinear activation eigenvectors and discussing the limitations of the leading oversmoothing measures, we propose the use of the rank of the features as a better measure of oversmoothing. We provide extensive experiments to validate the robustness of the effective rank against the classical measures. In addition, we have proved theoretically the decay of the rank of the features for message-passing GNNs.
Acknowledgement
KZ was supported by the EPSRC Centre for Doctoral Training in Mathematical Modelling, Analysis and Computation (MAC-MIGS) funded by the UK Engineering and Physical Sciences Research Council (grant EP/S023291/1), Heriot-Watt University and the University of Edinburgh. PD and FT are members of the Gruppo Nazionale Calcolo Scientifico - Istituto Nazionale di Alta Matematica (GNCS-INdAM). PD was supported by the MUR-PRO3 grant STANDS. DJH was supported by a Fellowship from the Leverhulme Trust. FT was partially funded by the PRIN-MUR project MOLE, code 2022ZK5ME7, and PRIN-PNRR project FIN4GEO within the European Union’s Next Generation EU framework, Mission 4, Component 2, CUP P2022BNB97.
@@ Appendix A Proofs of the main results
A.1 Proof of Theorem 3.2
@@
$\geq|\lambda_{1}|^{l}\|v_{1}^{\top}X^{(0)}W^{(0)}\dots W^{(l-1)}\|_{F}\Big{(}1-O\Big{(}{l\choose N}\Big{|}\frac{\lambda_{2}}{\lambda_{1}}\Big{|}^{l}\frac{\|X^{(0)}W^{(0)}\dots W^{(l-1)}\|_{F}}{\|v_{1}^{\top}X^{(0)}W^{(0)}\dots W^{(l-1)}\|_{F}}\Big{)}$
Now observe that under the randomness hypothesis from [11] and more generally from the Oseledets ergodic multiplicative theorem, we have that for almost any the limit $w\in\mathbb{R}^{d}$ $\lim_{l\rightarrow\infty}\frac{1}{l}\log\|w^{\top}W^{(0)}\dots W^{(l-1)}\|=c(\nu)$ exists and is equal to the maximal Lyapunov exponent of the system. In particular for any $w$ and $\epsilon>0$ there exists $l_{w,\epsilon}$ sufficiently large such that for any $l>l_{w,\epsilon}$
@@
$e^{l(c(\nu)-\epsilon)}\leq\|w^{\top}W^{(0)}\dots W^{(l-1)}\|<e^{l(c(\nu)+\epsilon)}\quad\forall l\geq l_{w,\epsilon}$ (14)
Now take as vector $w$ first the rows of $X^{(0)}$ and then the vector $v_{1}^{\top}X^{(0)}$, than almost surely for any $\epsilon$ there exists $l_{\epsilon}$ such that for any $l>l_{\epsilon}$
@@
($\left({l\choose N}\Big{|}\frac{\lambda_{2}}{\lambda_{1}}\Big{|}^{l}\frac{\|X^{(0)}W^{(0)}\dots W^{(l-1)}\|_{F}}{\|v_{1}^{\top}X^{(0)}W^{(0)}\dots W^{(l-1)}\|_{F}}\right)\leq\left({l\choose N}\Big{|}\frac{\lambda_{2}}{\lambda_{1}}\Big{|}^{l}e^{2l\epsilon}\right)$ (17)
@@
A.3 An illustrative example
Consider the stochastic nonnegative primitive matrix
$A:=\begin{pmatrix}0&1\\[4.30554pt] 1/2&1/2\end{pmatrix}.$ (19)
Since $A$ is row stochastic, the dominant eigenvector is given by the constant vector $u=(1,1)$. In particular given a vector $x=(x_{1},x_{2})\in\mathcal{K}$ we have that
$d_{H}(x,u)=\log\Big{(}\frac{\max_{i=1,2}\{x_{i}\}}{\min_{i=1,2}\{x_{i}\}}\Big{)}=\begin{cases}\log\big{(}x_{1}/x_{2}\big{)}\quad\text{if}\;x_{1}\geq x_{2}\\ \log\big{(}x_{2}/x_{1}\big{)}\quad\text{if}\;x_{2}\geq x_{1}\end{cases}.$ (20)
On the other hand $Ax=\big{(}x_{2}\,,\,(x_{1}+x_{2})/2\big{)}$, thus
$d_{H}(Ax,u)=\begin{cases}\log\big{(}(x_{1}+x_{2})/2x_{2}\big{)}\quad\text{if}\;x_{1}\geq x_{2}\\ \log\big{(}2x_{2}/(x_{1}+x_{2})\big{)}\quad\text{if}\;x_{2}\geq x_{1}\end{cases}.$ (21)
Now observe that the set of vectors $x$ such that $d_{H}(x,u)\leq C$ is given by the points $\{x\,|\;x_{1}\leq x_{2}\leq C^{*}x_{1}\}\cup\{x\,|\;x_{2}\leq x_{1}\leq C^{*}x_{2}\}$ with $C^{*}=e^{C}>1$. We want to prove the existence of some $\beta_{C}=1-\epsilon_{C}\in(0,1)$ such that $d_{H}(Ax,u)\leq\beta_{C}d_{H}(x,u)$ for any $x$ with $d_{H}(x,u)\leq C$.
To this end consider first the case of $x\in\Omega_{1}=\{x\,|\;x_{1}\leq x_{2}\leq C^{*}x_{1}\}$. In this case
$d_{H}(Ax,u)=\log\Big{(}\frac{2x_{2}}{x_{1}+x_{2}}\Big{)}\leq(1-\epsilon_{C})\log\Big{(}\frac{x_{2}}{x_{1}}\Big{)}=(1-\epsilon_{C})d_{H}(x,u)\qquad\forall x\in\Omega_{1}$ (22)
$\Longleftrightarrow\quad\epsilon_{C}\log\Big{(}\frac{x_{2}}{x_{1}}\Big{)}\leq\log\Big{(}\frac{x_{2}(x_{1}+x_{2})}{2x_{1}x_{2}}\Big{)}\qquad\forall x\in\Omega_{1}$
$\Longleftrightarrow\quad\epsilon_{C}\log(t)\leq\log\Big{(}\frac{1+t}{2}\Big{)}\qquad\forall t\in[1,C^{*}]$
$\Longleftrightarrow t^{\epsilon_{C}}\leq\frac{1+t}{2}\qquad\forall t\in[1,C^{*}]$
where we have written $t=x_{2}/x_{1}$. Second we consider the case of $x\in\Omega_{2}=\{x\,|\;x_{2}\leq x_{1}\leq C^{*}x_{2}\}$. In this case
$d_{H}(Ax,u)=\log\Big{(}\frac{x_{1}+x_{2}}{2x_{2}}\Big{)}\leq(1-\epsilon_{C})\log\Big{(}\frac{x_{1}}{x_{2}}\Big{)}=(1-\epsilon_{C})d_{H}(x,u)\qquad\forall x\in\Omega_{2}$ (23)
$\Longleftrightarrow\quad\epsilon_{C}\log\Big{(}\frac{x_{1}}{x_{2}}\Big{)}\leq\log\Big{(}\frac{2x_{1}x_{2}}{x_{2}(x_{1}+x_{2})}\Big{)}\qquad\forall x\in\Omega_{2}$
$\Longleftrightarrow\quad\epsilon_{C}\log(t)\leq\log\Big{(}\frac{2t}{t+1}\Big{)}\qquad\forall t\in[1,C^{*}]$
$\Longleftrightarrow t^{\epsilon_{C}}\leq\frac{2t}{t+1}\qquad\forall t\in[1,C^{*}]$
where we have written $t=x_{1}/x_{2}$. Both the functions $t\rightarrow(1+t)/2$ and $t\rightarrow 2t/(t+1)$ are monotonically increasing for $t\in[1,\infty)$ with derivative $1/2$ in $t=1$, thus for any $C<\infty$ there exists some $10<\epsilon_{C}<$ 1 such that both (22) and (22) are satisfied in the interval $t\in[1,C^{*}]$. This shows that $A$ is contractive with respect to $u$ even if $A$ is not strictly positive.
A.4 proof of Lemma 5.3
By the contraction properties of the matrix $A$ we know that if $\max_{i}d_{H}\big{(}(AX)_{:,i},u\big{)}\leq C$, then
@@ A.3 An illustrative example
for some $\beta<1$, where we have used $\lambda_{1}(A)>0$ and the scaling invariant property of the Hilbert distance.
Then note that, for any $i$, we can write $F(X)_{:,j}$ as follows
$\big{(}F(X)\big{)}_{:,j}=\sum_{j}W_{ij}(AX)_{:,j}\,.$ (25)
@@
Observe that if the claim holds we have conlcuded the proof, indeed by induction it can trivially be extended from $2$ to $d$ points yielding
$d_{H}\Big{(}F(X)_{:,j},u\Big{)}\leq\max_{i}d_{H}\Big{(}W_{ij}(AX)_{:,i},u\Big{)}\leq\max_{j}d_{H}\Big{(}(AX)_{:,j},u\Big{)}\leq\beta\max_{j}d_{H}\Big{(}X_{:,j},u\Big{)},$ (27)
@@
We only miss to prove the claim. To this end, exploiting the expression of the Hilbert distance we write
@@
A.5 proof of Theorem 5.4
To prove the theorem we start proving that, a continuous subhomogeneous and order-preserving function $\sigma$ with an eigenvector $u$ in the cone, is not nonexpansive in Hilbert distance with respect to $u$. Formally we claim that
@@
To prove it let $y\in\mathcal{K}$ and assume w.l.o.g. that $\|y\|_{1}=t>0$ and $\|u\|_{1}=1$, then
@@
$m(y/tu)\lambda_{t}tu\leq_{\mathcal{K}}\sigma\big{(}m(y/tu)tu\big{)}\leq_{\mathcal{K}}f(y)\leq_{\mathcal{K}}f\big{(}M(y/tu)tu\big{)}\leq_{\mathcal{K}}M(y/x_{c})\lambda_{t}tu.$ (31)
In particular we have $m(f(y)/tu)\geq\lambda_{t}m(y/tu)$ and $M(f(y)/tu)\leq\lambda_{t}M(y/tu)$. Finally the last inequalities and the scale invariance property of $d_{H}$ yield the thesis:
@@
Now using the claim above and Lemma 5.3, we can easily conclude that
$\lim_{l\rightarrow\infty}\max_{i}d_{H}(X_{:,i}^{(l)},u)=0.$ (33)
Indeed by the uniform contractivity of $\{A^{(l)}\}$, if $C=\max_{i}d_{H}(X^{(0)}_{:,i},u)$, then there exists some $\beta_{C}<1$ such that
$\max_{i}d_{H}(X^{(l)}_{:,i},u)\leq\beta_{C}^{l}\max_{i}d_{H}(X^{(0)}_{:,i},u).$ (34)
Finally to conclude we prove that, as a consequence of (33),
$\lim_{l\rightarrow\infty}\frac{\|(I-\mathcal{P})X^{(l)}\|_{F}}{\|\mathcal{P}X^{(l)}\|_{F}}=0.$ (35)
As a consequence of the inequality:
$1\leq\mathrm{NumRank}(X^{(l)})\leq 1+\frac{\|(I-\mathcal{P})X^{(l)}\|_{F}^{2}}{\|X^{(l)}\|_{2}^{2}}\leq 1+\frac{\|(I-\mathcal{P})X^{(l)}\|_{F}^{2}}{\|\mathcal{P}X^{(l)}\|_{F}^{2}},$ (36)
we have that (35) finally yields the thesis. To prove (35), we recall from Lemma 2.5.1 in [18] that for any $w$ such that $u^{\top}w=c$
@@
where $d_{T}(x,y)=\log(\max\{M(x/y),m(x/y)^{-1}\})$ and where since the dual cone of +$\mathbb{R}^{n}_{+}$ is +$\mathbb{R}^{n}_{+}$ itself, we are considering the norm induced by $u$ on the cone, i.e. $\|x\|_{u}=u^{\top}x$ for any $x$ in the cone. In practice the norm induce by $u$ $\|\cdot\|_{u}$ can be defines by the Minkowki functional of the set $\Omega=\text{ConvexHull}\{\{\Omega_{1}\}\cup\{-\Omega_{1}\}\}$ where $\Omega_{1}=\{x\in\mathcal{K}\,\;u^{\top}x\leq 1\}$
@@
the proof is concluded using (33) and observing that by the scale invariance property of the Hilbert distance $d_{H}(X^{(l)}_{:,i},\mathcal{P}X^{(l)}_{:,i})=d_{H}(X^{(l)}_{:,i},u)$, yielding:

### v4

@@ 6 Experiments
The vast majority of empirical studies in the literature that observe the decay of Dirichlet-like energy metrics are conducted over the layers of deep but untrained networks Wang et al. (2022); Rusch et al. (2022; 2023b); Wu et al. (2023); Roth (2024); Wang et al. (2025). Moreover, the measurements are done by looking at the layers of a single network, rather than different networks of increasing depth. We emphasize that this is an overly simplified and unrealistic setting. In this section, we perform an extensive numerical investigation on the behaviour of different oversmoothing metrics when measured on networks of different depths $l=2,\dots,L$, trained in isolation at different depths. Our analysis shows that GNN suffer from significant performance degradation after only few-layers, at which stage the convergence patterns of Dirichlet-like metrics are difficult to observe, while relaxed rank metric already show a significant decrease, well-correlating with the performance drop.
@@
  • The Dirichlet Energy $E_{\mathrm{Dir}}$ Cai and Wang (2020); Rusch et al. (2023a) and its variant $E_{\mathrm{Proj}}$ Wu et al. (2023). Both are discussed in Section˜3.1, see in particular (1) and (2).
  • Normalized versions of Dirichlet energy and its variant, $E_{\mathrm{Dir}}(X)/\|X\|_{F}^{2}$ and $E_{\mathrm{Proj}}(X)/\|X\|_{F}$. Indeed, from our previous discussion, a robust oversmoothing measure should be scale invariant with respect to the features. Metrics with global normalization like the ones we consider here have also been proposed in Di Giovanni et al. (2023); Roth and Liebig (2023); Maskey et al. (2023).
  • The Mean Average Distance (MAD) Chen et al. (2020)
@@
  • Relaxed rank metrics: We consider the Numerical Rank and Effective Rank. Both are discussed in Section˜5. We point out that from our theoretical investigation, in particular from (4), the numerical rank decays to $1$ faster than the decay of the normalized $E_{\mathrm{Proj}}$ energy to zero. This further supports the use of the Numerical Rank as an improved measure of oversmoothing.
[…]
In Table˜1 and Figure˜2, we train GNNs of a fixed hidden dimension equal to 32 on homophilic, heterophilic and large-scale datasets in their default splits. We follow the standard setups of GCN and GAT as stated in Sections˜2.1 and 2.2, and use homogeneous LeakyReLU (LReLU) as the activation function. For each configuration, GNNs of eight different depths ranging from 2 to 24 are trained. The oversmoothing metric and accuracy results are averaged over 10 separately trained GNNs. All GNNs are trained with NAdam Optimizer and a constant learning rate of 0.01. The oversmoothing metrics are computed at the last hidden layer before the output layer. In Figure˜2 and in Appendix˜G, we plot the behaviour of the different oversmoothing measures, the norm of the features, and the accuracy of the trained GNNs with increasing depth. These figures clearly show that the network suffers a significant drop in accuracy, which is not matched by any visible change in standard oversmoothing metrics. By contrast, the rank of the feature representations decreases drastically, following closely the behaviour of the network’s accuracy. These findings are further supported by the results shown in Table˜1, where we compute the Pearson correlation coefficient between the logarithm of every measure and the classification accuracy of every GNN model. The use of a logarithmic transformation is based on the understanding that oversmoothing grows exponentially with the length of the network.
Extensions of these results are provided in Appendices˜E and F. In Appendix˜D, we perform an asymptotic ablation study on very deep (300-layer) synthetic networks with randomly sampled, untrained weights. This study serves to validate our theoretical findings on the convergence of relaxed rank metrics and to demonstrate that such untrained settings offer little insight into the ability of existing metrics to quantify oversmoothing in realistic, trained networks.
@@ 7 Conclusion
In this paper, we have discussed the problem of quantifying oversmoothing in message-passing GNNs. After simplifying the existing theoretical analysis using nonlinear activation eigenvectors and discussing the limitations of the leading oversmoothing measures, we propose the use of the rank of the features as a better measure of oversmoothing. We provide extensive experiments to validate the robustness of the effective rank against the classical measures. In addition, we have analysed theoretically the decay of the rank of the features for message-passing GNNs. Notably, our work opens new possible research directions that investigate the use of rank-based metrics in novel architectures that can alleviate the oversmoothing phenomenon. Indeed, Dirichlet-like metrics have been employed not only to measure oversmoothing but also to design new architectures able to mitigate this phenomenon, e.g. by considering residual architectures where the magnitude of the residual is regulated according to the Dirichlet energy of neighboring nodes Rusch et al. (2023b); Jin and Zhu (2024). In light of our results and the superiority of rank-based metrics in capturing oversmoothing, it is natural to wonder whether the use of rank-based metrics in place of Dirichlet-like ones could better mitigate oversmoothing in similar kind of architectures. We consider this as a promising future research line.
8 Acknowledgement
KZ was supported by the EPSRC Centre for Doctoral Training in Mathematical Modelling, Analysis and Computation (MAC-MIGS) funded by the UK Engineering and Physical Sciences Research Council (grant EP/S023291/1), Heriot-Watt University and the University of Edinburgh. PD and FT are members of the Gruppo Nazionale Calcolo Scientifico - Istituto Nazionale di Alta Matematica (GNCS-INdAM). PD was supported by the INdAM-GNCS project “NLA4ML—Numerical Linear Algebra Techniques for Machine Learning”. FT is partially funded by the PRIN-MUR project MOLE (code 2022ZK5ME7) and the PRIN-PNRR project FIN4GEO (code P2022BNB97). DJH was supported by the Advanced Grant “Numerical Analysis for Stable AI” 101198795 from the European Research Council.
@@ Appendix A Proofs of the main results
A.1 Equivalence of $E_{\mathrm{Proj}}$ and $E_{\mathrm{Dir}}$
Lemma A.1.
Assume the graph $\mathcal{G}$ to be connected and the eigenvector $u$ to be strictly positive $u_{i}>0$ for all $i$ and such that $\|u\|_{2}=1$. Then there exist $C_{1}>0$ and $C_{2}>0$ such that
$C_{1}E_{\mathrm{Dir}}(X)\leq E_{\mathrm{Proj}}(X)\leq C_{2}E_{\mathrm{Dir}}(X)\qquad\forall X\in\mathbb{R}^{N\times d}.$
Proof.
First we show that the Dirichlet energy is equivalent to the modified Dirichlet energy where all the pairs of nodes $(i,j)$ are considered
$E_{\mathrm{Dir}}(X)=\sum_{(i,j)\in\mathcal{E}}\left\|\frac{X_{i,:}}{u_{i}}-\frac{X_{j,:}}{u_{j}}\right\|^{2}_{2}\leq\sum_{i\in\mathcal{V}}\sum_{j\in\mathcal{V}}\left\|\frac{X_{i,:}}{u_{i}}-\frac{X_{j,:}}{u_{j}}\right\|^{2}_{2}=\tilde{E_{\mathrm{Dir}}}(X).$ (7)
Second observe that if $i,j\in\mathcal{V}$, since the graph is connected there exists some path $i_{1}=1,i_{2},\dots,i_{n+1}=j$ such that $(i_{k},i_{k+1})\in\mathcal{E}$ for all $k$. Thus, using the traingular inequality we get
$\left\|\frac{X_{i,:}}{u_{i}}-\frac{X_{j,:}}{u_{j}}\right\|^{2}_{2}\leq n\sum_{k=1}^{n}\left\|\frac{X_{i_{k},:}}{u_{i_{k}}}-\frac{X_{i_{k+1},:}}{u_{i_{k+1}}}\right\|^{2}_{2}\leq nE_{\mathrm{Dir}}(X).$ (8)
repeating the same argument for any pair of nodes $(i,j)$ we observe that for some constant $C>1$
$\tilde{E_{\mathrm{Dir}}}(X)\leq\widetilde{C}E_{\mathrm{Dir}}(X).$ (9)
So, to prove the equivalence of $E_{\mathrm{Dir}}$ and $E_{\mathrm{Proj}}$ it is sufficient to prove the equivalence between $E_{\mathrm{Proj}}$ and $\tilde{E_{\mathrm{Dir}}}$.
If we make explicit the expression of the norms in $\tilde{E_{\mathrm{Dir}}}$ we get
$\tilde{E_{\mathrm{Dir}}}(X)=$ $\sum_{i\in\mathcal{V}}\sum_{j\in\mathcal{V}}\sum_{k=1}^{d}\left|\frac{X_{i,k}}{u_{i}}-\frac{X_{j,k}}{u_{j}}\right|^{2}=\sum_{k=1}^{d}\sum_{i\in\mathcal{V}}\sum_{j\in\mathcal{V}}\left|\frac{X_{i,k}}{u_{i}}-\frac{X_{j,k}}{u_{j}}\pm u^{T}X_{:,k}\right|^{2}$ (10)
$\leq\sum_{k=1}^{d}\sum_{i\in\mathcal{V}}\sum_{j\in\mathcal{V}}\frac{2}{u_{i}}\left|X_{i,k}-(u^{T}X_{:,k})u_{i}\right|^{2}+\frac{2}{u_{j}}\left|X_{j,k}-(u^{T}X_{:,k})u_{j}\right|^{2}$
$\leq\frac{4|\mathcal{V}|}{\min_{i}\{u_{i}\}}\sum_{k=1}^{d}\left\|X_{:,k}-uu^{T}X_{:,k}\right\|^{2}=\frac{4|\mathcal{V}|}{\min_{i}\{u_{i}\}}E_{\mathrm{Proj}}(X).$
To prove the opposite observe the following
$E_{\mathrm{Proj}}(X)=$ $\sum_{k=1}^{d}\sum_{i\in\mathcal{V}}u_{i}\left|\frac{X_{i,k}}{u_{i}}-(u^{T}X_{:,k})\right|^{2}=\sum_{k=1}^{d}\sum_{i\in\mathcal{V}}u_{i}\left|\frac{X_{i,k}}{u_{i}}-\left(\sum_{h}u_{h}^{2}\frac{X_{h,k}}{u_{h}}\right)\right|^{2}$ (11)
$\leq\max_{j}\{u_{j}\}\sum_{k=1}^{d}\sum_{i\in\mathcal{V}}\max_{j}\left|\frac{X_{i,k}}{u_{i}}-\frac{X_{j,k}}{u_{j}}\right|^{2}$
$\leq\max_{j}\{u_{j}\}\sum_{k=1}^{d}\sum_{i\in\mathcal{V}}\sum_{j\in\mathcal{V}}\left|\frac{X_{i,k}}{u_{i}}-\frac{X_{j,k}}{u_{j}}\right|^{2}=\tilde{E_{\mathrm{Dir}}}(X)$
where in the first inequality we have used that, since $\|u\|_{2}=1$, $\sum_{h}u_{h}^{2}\frac{X_{h,k}}{u_{h}}$ is a convex combination of $\left\{\frac{X_{h,k}}{u_{h}}\right\}_{h}$. In particular, the last inequality concludes the proof.
A.2 Proof of Theorem˜3.2
@@ A.1 Equivalence of $E_{\mathrm{Proj}}$ and $E_{\mathrm{Dir}}$
$\geq|\lambda_{1}|^{l}\|v_{1}^{\top}X^{(0)}W^{(0)}\dots W^{(l-1)}\|_{F}\Big(1-O\Big({l\choose N}\Big|\frac{\lambda_{2}}{\lambda_{1}}\Big|^{l}\frac{\|X^{(0)}W^{(0)}\dots W^{(l-1)}\|_{F}}{\|v_{1}^{\top}X^{(0)}W^{(0)}\dots W^{(l-1)}\|_{F}}\Big)\$
Now observe that under the randomness hypothesis from furstenbergRandomMatrixProducts1983 and more generally from the Oseledets ergodic multiplicative theorem, we have that for almost any the limit $w\in\mathbb{R}^{d}$ $\lim_{l\rightarrow\infty}\frac{1}{l}\log\|w^{\top}W^{(0)}\dots W^{(l-1)}\|=c(\nu)$ exists and is equal to the maximal Lyapunov exponent of the system. In particular for any $w$ and $\epsilon>0$ there exists $l_{w,\epsilon}$ sufficiently large such that for any $l>l_{w,\epsilon}$
@@
$e^{l(c(\nu)-\epsilon)}\leq\|w^{\top}W^{(0)}\dots W^{(l-1)}\|<e^{l(c(\nu)+\epsilon)}\quad\forall l\geq l_{w,\epsilon}.$ (20)
Now take as vector $w$ first the rows of $X^{(0)}$ and then the vector $v_{1}^{\top}X^{(0)}$, then almost surely for any $\epsilon$ there exists $l_{\epsilon}$ such that for any $l>l_{\epsilon}$
@@
$\left({l\choose N}\Big|\frac{\lambda_{2}}{\lambda_{1}}\Big|^{l}\frac{\|X^{(0)}W^{(0)}\dots W^{(l-1)}\|_{F}}{\|v_{1}^{\top}X^{(0)}W^{(0)}\dots W^{(l-1)}\|_{F}}\right)\leq\left({l\choose N}\Big|\frac{\lambda_{2}}{\lambda_{1}}\Big|^{l}e^{2l\epsilon}\right)$ (23)
@@
A.4 Proof of Lemma˜5.2
Since $A$ is nonnegative, then from Perron Frobenius theory lemmensNonlinearPerronFrobeniusTheory2012, we know that
@@
for some $\beta\leq 1$, where we have used $\lambda_{1}(A)>0$ and the scaling invariant property of the Hilbert distance. In particular if $A$ is contractive in Hilbert distance $\beta<1$.
Then note that, for any $i$, we can write $(AXW)_{:,j}$ as follows
$\big(AXW\big)_{:,j}=\sum_{j}W_{ij}(AX)_{:,j}\,.$ (26)
@@
Observe that if the claim holds, by induction it can trivially be extended from $2$ to $d$ points yielding
$d_{H}\Big((AXW)_{:,j},u\Big)\leq\max_{i}d_{H}\Big(W_{ij}(AX)_{:,i},u\Big)\leq\max_{j}d_{H}\Big((AX)_{:,j},u\Big)\leq\beta\max_{j}d_{H}\Big(X_{:,j},u\Big),$ (28)
@@
We miss to prove the claim. To this end, exploiting the expression of the Hilbert distance we write
@@
Next we prove that, a continuous subhomogeneous and order-preserving function $\sigma$ with eigenvector $u$ in the cone, is not nonexpansive in Hilbert distance with respect to $u$. Formally we claim that
@@
To prove it, let $y\in\mathcal{K}$ and assume w.l.o.g. that $\|y\|_{1}=t>0$ and $\|u\|_{1}=1$, then
@@
$m(y/tu)\lambda_{t}tu\leq_{\mathcal{K}}\sigma\big(m(y/tu)tu\big)\leq_{\mathcal{K}}f(y)\leq_{\mathcal{K}}f\big(M(y/tu)tu\big)\leq_{\mathcal{K}}M(y/x_{c})\lambda_{t}tu,$ (32)
where the inequalities are strict if $\sigma$ is strictly subhomogeneous. In particular we have $m(f(y)/tu)\geq\lambda_{t}m(y/tu)$ and $M(f(y)/tu)\leq\lambda_{t}M(y/tu)$. Finally the last inequalities and the scale invariance property of $d_{H}$ yield the thesis:
@@
with the inequality that is strict if $\sigma$ is strictly subhomogeneous.
Then the thesis of the Lemma follows by applying (30) to (28).
A.5 proof of Theorem˜5.3
We will prove that, as a consequence of the hypothesis $\lim_{l\rightarrow\infty}\max_{i}d_{H}(X_{:,i}^{(l)},u)=0$,
$\lim_{l\rightarrow\infty}\frac{\|(I-\mathcal{P})X^{(l)}\|_{F}}{\|\mathcal{P}X^{(l)}\|_{F}}=0,$ (34)
where $\mathcal{P}=uu^{T}/\|u\|_{2}^{2}$. Indeed (34) is equivalent to proving that the numerical rank goes to 1 as $l$ goes to $\infty$:
$1\leq\mathrm{NumRank}(X^{(l)})\leq 1+\frac{\|(I-\mathcal{P})X^{(l)}\|_{F}^{2}}{\|X^{(l)}\|_{2}^{2}}\leq 1+\frac{\|(I-\mathcal{P})X^{(l)}\|_{F}^{2}}{\|\mathcal{P}X^{(l)}\|_{F}^{2}}.$ (35)
To prove (34), we recall from Lemma 2.5.1 in lemmensNonlinearPerronFrobeniusTheory2012 that for any $w$ such that $u^{\top}w=c$
@@
where $d_{T}(x,y)=\log(\max\{M(x/y),m(x/y)^{-1}\})$ and where since the dual cone of $\mathbb{R}^{n}_{+}$ is $\mathbb{R}^{n}_{+}$ itself, we are considering the norm induced by $u$ on the cone, i.e. $\|x\|_{u}=u^{\top}x$ for any $x$ in the cone. In practice the norm induce by $u$ $\|\cdot\|_{u}$ can be defines by the Minkowki functional of the set $\Omega=\text{ConvexHull}\{\{\Omega_{1}\}\cup\{-\Omega_{1}\}\}$ where $\Omega_{1}=\{x\in\mathcal{K}\,\;u^{\top}x\leq 1\}$
@@
the proof is concluded using the hypothesis and observing that by the scale invariance property of the Hilbert distance $d_{H}(X^{(l)}_{:,i},\mathcal{P}X^{(l)}_{:,i})=d_{H}(X^{(l)}_{:,i},u)$, yielding: